OpenStudy (anonymous):

We want to find how much charge is on the electrons in a nickel coin. Follow this method. A nickel coin has a mass of about 4.4 g. Find the number of atoms in a nickel coin. Each mole (6.02 × 1023 atoms) has a mass of about 58.1 g. Answer in units of atoms

OpenStudy (richyw):

atoms are not a unit of charge. So I don't know what the first sentence is about. but to get the number or atoms in your coin just use \[4.4\text{g}\frac{1\text{mol}}{58.2\text{g}}\frac{6.022\times 10^23 \text{atoms}}{\text{mol}\]

OpenStudy (richyw):

\[4.4\text{g}\frac{1\text{mol}}{58.2\text{g}}\frac{6.022\times 10^23 \text{atoms}}{\text{mol}\]

OpenStudy (richyw):

$$4.4\text{g}\frac{1\text{mol}}{58.2\text{g}}\frac{6.022\times 10^23 \text{atoms}}{\text{mol}$$

OpenStudy (richyw):

what happened to the LaTeX?